# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ A.3 Proof of Theorem 2
For the newsvendor problem with the cost function given by $\ell(\xi,z)=wz+b(\xi-z)_{+}+h(z-\xi)_{+}$, DRO with type-$\infty$ Wasserstein ambiguity set is equivalent to
$\begin{array}[]{ll}\mathrm{minimize}&\displaystyle\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ {\rm subject\mbox{ }\rm to}&\displaystyle(\hat{\xi}_{i}+\epsilon)b-z(w-b)\leq t_{i}\,,\;\forall i\in[N]\\ &(\epsilon-\hat{\xi}_{i})h+z(w+h)\leq t_{i}\,,\;\forall i\in[N].\end{array}$ (33)
Proof.
For the newsvendor problem, the cost function is given by $\ell(\xi,z)=wz+b(\xi-z)_{+}+h(z-\xi)_{+}$. (23b) is equivalent to
$\mathop{\rm min}_{{{\boldsymbol{t}}},{{\boldsymbol{z}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha t_{i})\right)$ (34a)
subject ${\rm subject\mbox{ }\rm to}\quad$ $t_{i}\geq wz+b\xi-bz$ $\forall\xi_{i}\in\Xi^{i}_{\epsilon},\;\forall i\in[N]$ (34b)
$t_{i}\geq wz+hz-h\xi$ $\forall\xi_{i}\in\Xi^{i}_{\epsilon},\;\forall i\in[N],$ (34c)
where $\Xi^{i}_{\epsilon}=\{\xi:\|\xi-\hat{\xi}_{i}\|\leq\epsilon\}$. Define $\ell^{+}(\xi,z):=wz+b\xi-bz$ and $\ell^{-}(\xi,z)=wz+hz-h\xi$ for which the partial concave conjugates are given by:
$\ell^{+}_{*}(z,\varphi^{+})=\left\{\begin{array}[]{cl}z(b-w)&\text{if }\varphi^{+}=b\\ -\infty&\text{otherwise},\end{array}\right.$ (37)
$\ell^{-}_{*}(z,\varphi^{-})=\left\{\begin{array}[]{cl}-z(w+h)&\text{if }\varphi^{-}=-h\\ -\infty&\text{otherwise}.\end{array}\right.$ (40)
Thus, problem (27) reduces to (33).
@@
For the regression problem with loss function given by $\ell({{\boldsymbol{\xi}}},{{\boldsymbol{z}}})=|{{\boldsymbol{\xi}}}_{n+1}-{{\boldsymbol{\xi}}}_{1:n}^{T}{{\boldsymbol{z}}}|$, DRO with type-$\infty$ Wasserstein ambiguity set is equivalent to
@@
$\mathop{\rm min}_{{{\boldsymbol{z}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha|\hat{{{\boldsymbol{\xi}}}}_{n+1}^{i}-\hat{{{\boldsymbol{\xi}}}}_{1:n}^{i^{T}}{{\boldsymbol{z}}}|)\right)+\epsilon\|[-1\;{{\boldsymbol{z}}}^{T}]^{T}\|_{*}$ (41)
Proof.
@@
For the regression problem, the cost function is given by $\ell({{\boldsymbol{\xi}}},{{\boldsymbol{z}}})=|{{\boldsymbol{\xi}}}_{n+1}-{{\boldsymbol{\xi}}}_{1:n}^{T}{{\boldsymbol{z}}}|$. (23b) is equivalent to
$\mathop{\rm min}_{{{\boldsymbol{t}}},{{\boldsymbol{z}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha t_{i})\right)$ (42a)
subject ${\rm subject\mbox{ }\rm to}\quad$ $t_{i}\geq{{\boldsymbol{\xi}}}_{n+1}-{{\boldsymbol{\xi}}}_{1:n}^{T}{{\boldsymbol{z}}}$ $\forall\hat{{{\boldsymbol{\xi}}}}_{i}\in\Xi^{i}_{\epsilon},\;\forall i\in[N]$ (42b)
$t_{i}\geq{{\boldsymbol{\xi}}}_{1:n}^{T}{{\boldsymbol{z}}}-{{\boldsymbol{\xi}}}_{n+1}$ $\forall\hat{{{\boldsymbol{\xi}}}}_{i}\in\Xi^{i}_{\epsilon},\;\forall i\in[N],$ (42c)
where $\Xi^{i}_{\epsilon}=\{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon\}$. Define $\ell^{+}({{\boldsymbol{\xi}}},{{\boldsymbol{z}}}):={{\boldsymbol{\xi}}}_{n+1}-{{\boldsymbol{\xi}}}_{1:n}^{T}{{\boldsymbol{z}}}$ and $\ell^{-}({{\boldsymbol{\xi}}},{{\boldsymbol{z}}}):={{\boldsymbol{\xi}}}_{1:n}^{T}{{\boldsymbol{z}}}-{{\boldsymbol{\xi}}}_{n+1}$. Then, the partial concave conjugates of $\ell^{+}({{\boldsymbol{\xi}}},{{\boldsymbol{z}}})$ and $\ell^{-}({{\boldsymbol{\xi}}},{{\boldsymbol{z}}})$ are given by
$\ell^{+}_{*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}^{+})=\left\{\begin{array}[]{cl}0&\text{if }{{\boldsymbol{\varphi}}}^{+}_{n+1}=1\,\&\,{{\boldsymbol{\varphi}}}^{+}_{1:n}=-{{\boldsymbol{z}}}\\ -\infty&\text{otherwise},\end{array}\right.$ (45)
$\ell^{-}_{*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}^{-})=\left\{\begin{array}[]{cl}0&\text{if }{{\boldsymbol{\varphi}}}^{-}_{n+1}=-1\,\&\,{{\boldsymbol{\varphi}}}^{-}_{1:n}={{\boldsymbol{z}}}\\ -\infty&\text{otherwise}.\end{array}\right.$ (48)
Hence, the problem (27) reduces to
$\mathop{\rm min}_{{{\boldsymbol{t}}},{{\boldsymbol{z}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha t_{i})\right)$ (49a)
subject ${\rm subject\mbox{ }\rm to}\quad$ $t_{i}\geq\hat{{{\boldsymbol{\xi}}}}_{n+1}^{i}-\hat{{{\boldsymbol{\xi}}}}_{1:n}^{i^{T}}{{\boldsymbol{z}}}+\epsilon\|[1\;-{{\boldsymbol{z}}}^{T}]^{T}\|_{*}$ $\forall i\in[N]$ (49b)
$t_{i}\geq\hat{{{\boldsymbol{\xi}}}}_{1:n}^{i^{T}}{{\boldsymbol{z}}}-\hat{{{\boldsymbol{\xi}}}}_{n+1}^{i}+\epsilon\|[-1\;{{\boldsymbol{z}}}^{T}]^{T}\|_{*}$ $\forall i\in[N].$ (49c)
And finally,
@@
$\mathop{\rm min}_{{{\boldsymbol{z}}}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\exp(\alpha|\hat{{{\boldsymbol{\xi}}}}_{n+1}^{i}-\hat{{{\boldsymbol{\xi}}}}_{1:n}^{i^{T}}{{\boldsymbol{z}}}|)\right)+\epsilon\|[-1\;{{\boldsymbol{z}}}^{T}]^{T}\|_{*}.$ (50)
Appendix B Additional results
@@
Suppose that $\mathcal{Z}$ is compact, ${{\boldsymbol{\xi}}}_{1},{{\boldsymbol{\xi}}}_{2},\cdots,{{\boldsymbol{\xi}}}_{N}$ are i.i.d with distribution ${\mathbb{P}}$, and the loss function $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is Lipschitz continuous in ${{\boldsymbol{z}}}$ with Lipschitz constant $L({{\boldsymbol{\xi}}})$. Also, suppose that there exists a $p\geq 2$ for which ${\mathbb{E}}_{\mathbb{P}}\left[\left(L({{\boldsymbol{\xi}}})\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\right)^{p}\right]<\infty$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$. Then, it can be shown that for any $\gamma>0$, there exists a constant $A>0$, that is independent of $N$ for which,
@@
${\mathbb{P}}\left(\lvert\hat{\rho}_{\text{SAA}}-\rho^{*}\rvert\geq\gamma\right)\leq\frac{A}{N^{\frac{p}{2}}(\alpha\gamma\exp(\alpha\rho^{*}))^{p}}.$ (51)
for sufficiently large $N$.
Proof. From the local lipschitz continuity of $\ell$ in ${{\boldsymbol{z}}}$ for ${{\boldsymbol{\xi}}}\in\Xi$, we have:
$|\ell({{\boldsymbol{z}}}_{1},{{\boldsymbol{\xi}}})-\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}})|\leq L({{\boldsymbol{\xi}}})\|{{\boldsymbol{z}}}_{1}-{{\boldsymbol{z}}}_{2}\|,$ (52)
where $L({{\boldsymbol{\xi}}})$ is the lipschitz constant. From the Mean Value Theorem, we obtain:
$\lvert\exp(\alpha\ell({{\boldsymbol{z}}}_{1},{{\boldsymbol{\xi}}}))-\exp(\alpha\ell({{\boldsymbol{z}}}_{2},{{\boldsymbol{\xi}}}))\rvert$ $\ =\ \lvert\left(\left.\nabla_{{{\boldsymbol{z}}}}\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\right|_{{{\boldsymbol{z}}}=z_{\theta}}\right)^{\top}\left({{\boldsymbol{z}}}_{1}-{{\boldsymbol{z}}}_{2}\right)\rvert$
$\ =\alpha\exp(\alpha\ell({{\boldsymbol{z}}}_{\theta},{{\boldsymbol{\xi}}}))\lvert\left(\nabla_{{{\boldsymbol{z}}}}\ell({{\boldsymbol{z}}}_{\theta},{{\boldsymbol{\xi}}})\right)^{\top}\left({{\boldsymbol{z}}}_{1}-{{\boldsymbol{z}}}_{2}\right)\rvert$
$\ \leq\ \alpha\exp(\alpha\ell({{\boldsymbol{z}}}_{\theta},{{\boldsymbol{\xi}}}))\|\nabla_{{{\boldsymbol{z}}}}\ell({{\boldsymbol{z}}}_{\theta},{{\boldsymbol{\xi}}})\|\|{{\boldsymbol{z}}}_{1}-{{\boldsymbol{z}}}_{2}\|$
$\ \leq\ \alpha L({{\boldsymbol{\xi}}})\exp(\alpha\ell(z_{\theta},{{\boldsymbol{\xi}}}))\lVert{{\boldsymbol{z}}}_{1}-{{\boldsymbol{z}}}_{2}\rVert$
where ${{\boldsymbol{z}}}_{\theta}=\theta{{\boldsymbol{z}}}_{1}+(1-\theta){{\boldsymbol{z}}}_{2},\theta\in(0,1)$ and the first inequality follows from the Cauchy-Schwarz inequality and the second from (52). Since $\mathcal{Z}$ is a compact set, and there exists some $p\geq 2$ for which ${\mathbb{E}}[L({{\boldsymbol{\xi}}})\exp(p\alpha\ell({{\boldsymbol{z}}}_{1},{{\boldsymbol{\xi}}}))]<\infty$, it follows from Theorem 3.2 in Jiang et al., (2020) that for any $\gamma>0$, there exists $A>0$, independent of $N$, such that
${\mathbb{P}}\left(\mathop{\rm sup}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\left|{\mathbb{E}}_{{\mathbb{{\mathbb{P}}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))-{\mathbb{E}}_{{\mathbb{P}}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))\right|\geq\gamma\right)\leq\frac{A}{N^{\frac{p}{2}}\gamma^{p}}.$ (53)
for sufficiently large $N$. Denote by $\hat{u}_{N}({{\boldsymbol{z}}}):={\mathbb{E}}_{{\mathbb{{\mathbb{P}}}}_{N}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))$ and by $u({{\boldsymbol{z}}}):={\mathbb{E}}_{{\mathbb{P}}}(\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})))$. From the concavity of $\log$, it follows that:
$\log(\hat{u}_{N}({{\boldsymbol{z}}}))-\log(u({{\boldsymbol{z}}}))$ $\ \leq\ \frac{1}{u({{\boldsymbol{z}}})}\left(\hat{u}_{N}({{\boldsymbol{z}}})-u({{\boldsymbol{z}}})\right)\;\forall z\in\mathcal{Z}$

### v3

@@ A.8 Additional results
The DRO newsvendor problem with the cost function given by $\ell(\xi,z)=wz+b(\xi-z)_{+}+h(z-\xi)_{+}$ and type-$\infty$ Wasserstein ambiguity set is equivalent to:
$\begin{array}[]{lll}\mathrm{minimize}&\displaystyle\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ {\rm subject\mbox{ }\rm to}&\displaystyle(\hat{\xi}_{i}+\epsilon)b-z(w-b)\leq t_{i}&\forall i\in[N]\\ &(\epsilon-\hat{\xi}_{i})h+z(w+h)\leq t_{i}&\forall i\in[N].\end{array}$ (32)
Proof. For the newsvendor problem, the cost function is given by $\ell(z,\xi)=wz+b(\xi-z)_{+}+h(z-\xi)_{+}$. Problem (30) is equivalent to:
$\begin{array}[]{lll}\mathop{\rm min}&\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ \text{s.t.}&\boldsymbol{t}\in\mathbb{R}^{N},\,z\in\mathbb{R}\\ &t_{i}\geq wz+b\xi-bz&\forall\xi\in\Xi^{i}_{\epsilon},\,\forall i\in[N]\\ &t_{i}\geq wz+hz-h\xi&\forall\xi\in\Xi^{i}_{\epsilon},\,\forall i\in[N],\end{array}$ (33)
where $\Xi^{i}_{\epsilon}=\{\xi:\|\xi-\hat{\xi}_{i}\|\leq\epsilon\}$. Define $\ell^{+}(z,\xi):=wz+b\xi-bz$ and $\ell^{-}(z,\xi)=wz+hz-h\xi$ for which the partial concave conjugates are given by:
$\begin{array}[]{lll}&\ell^{+}_{*}(z,\varphi^{+})=\left\{\begin{array}[]{cl}z(b-w)&\text{if }\varphi^{+}=b\\ -\infty&\text{otherwise},\end{array}\right.\\ \\ &\ell^{-}_{*}(z,\varphi^{-})=\left\{\begin{array}[]{cl}-z(w+h)&\text{if }\varphi^{-}=-h\\ -\infty&\text{otherwise}.\end{array}\right.\end{array}$
Thus, from Theorem 5, problem (30) reduces to (32).
@@
The DRO regression problem with loss function $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})=|{{\boldsymbol{\xi}}}_{d+1}-{{\boldsymbol{\xi}}}_{1:d}^{\top}{{\boldsymbol{z}}}|$, and type-$\infty$ Wasserstein ambiguity set is equivalent to:
@@
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathbb{R}^{d}}\quad$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha|\hat{{{\boldsymbol{\xi}}}}_{i,d+1}-\hat{{{\boldsymbol{\xi}}}}_{i,1:d}^{\top}{{\boldsymbol{z}}}|)\right)+\epsilon\|[-1\;{{\boldsymbol{z}}}^{\top}]^{\top}\|_{*}.$ (34)
@@
Proof. For the regression problem, the cost function is given by $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})=|{{\boldsymbol{\xi}}}_{d+1}-{{\boldsymbol{\xi}}}_{1:d}^{T}{{\boldsymbol{z}}}|$. Problem (30) is equivalent to:
@@
$\begin{array}[]{ll}\mathop{\rm min}&\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ \text{s.t.}&\boldsymbol{t}\in\mathbb{R}^{N},\,\boldsymbol{z}\in\mathbb{R}^{d}\\ &t_{i}\geq{{\boldsymbol{\xi}}}_{d+1}-{{\boldsymbol{\xi}}}_{1:d}^{\top}{{\boldsymbol{z}}}\quad\forall{{\boldsymbol{\xi}}}\in\Xi^{i}_{\epsilon},\;\forall i\in[N]\\ &t_{i}\geq{{\boldsymbol{\xi}}}_{1:d}^{\top}{{\boldsymbol{z}}}-{{\boldsymbol{\xi}}}_{d+1}\quad\forall{{\boldsymbol{\xi}}}\in\Xi^{i}_{\epsilon},\;\forall i\in[N],\end{array}$ (35)
where $\Xi^{i}_{\epsilon}=\{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon\}$. Define $\ell^{+}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}):={{\boldsymbol{\xi}}}_{d+1}-{{\boldsymbol{\xi}}}_{1:d}^{\top}{{\boldsymbol{z}}}$ and $\ell^{-}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}):={{\boldsymbol{\xi}}}_{1:d}^{\top}{{\boldsymbol{z}}}-{{\boldsymbol{\xi}}}_{d+1}$. Then, the partial concave conjugates of $\ell^{+}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ and $\ell^{-}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ are given by:
$\ell^{+}_{*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}^{+})=\left\{\begin{array}[]{cl}0&\text{if }{{\boldsymbol{\varphi}}}^{+}_{d+1}=1\,\text{ and }\,{{\boldsymbol{\varphi}}}^{+}_{1:d}=-{{\boldsymbol{z}}}\\ -\infty&\text{otherwise},\end{array}\right.$ (36c)
$\ell^{-}_{*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}^{-})=\left\{\begin{array}[]{cl}0&\text{if }{{\boldsymbol{\varphi}}}^{-}_{d+1}=-1\,\text{ and }\,{{\boldsymbol{\varphi}}}^{-}_{1:d}={{\boldsymbol{z}}}\\ -\infty&\text{otherwise}.\end{array}\right.$ (36f)
Substituting (36) in problem (30) reduces to:
$\begin{array}[]{lll}\mathop{\rm min}&\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ \text{s.t.}&\boldsymbol{t}\in\mathbb{R}^{N},\,\boldsymbol{z}\in\mathbb{R}^{d}\\ &t_{i}\geq\hat{{{\boldsymbol{\xi}}}}_{i,d+1}-\hat{{{\boldsymbol{\xi}}}}_{i,1:d}^{\top}{{\boldsymbol{z}}}+\epsilon\|[1\;-{{\boldsymbol{z}}}^{\top}]^{\top}\|_{*}&\forall i\in[N]\\ &t_{i}\geq\hat{{{\boldsymbol{\xi}}}}_{i,1:d}^{\top}{{\boldsymbol{z}}}-\hat{{{\boldsymbol{\xi}}}}_{i,d+1}+\epsilon\|[-1\;{{\boldsymbol{z}}}^{\top}]^{\top}\|_{*}&\forall i\in[N].\end{array}$
Finally:
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathbb{R}^{d}}\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha|\hat{{{\boldsymbol{\xi}}}}_{i,d+1}-\hat{{{\boldsymbol{\xi}}}}_{i,1:d}^{\top}{{\boldsymbol{z}}}|)\right)+\epsilon\|[-1\;{{\boldsymbol{z}}}^{\top}]^{\top}\|_{*}.$
@@
The following inequalities follow from the properties of the logarithm function:
$\begin{array}[]{ll}\log(1+\epsilon)\leq\epsilon&\text{ if }\epsilon\geq 0,\\ \log(1-\epsilon)\geq-\epsilon/(1-1/e)&\text{ if }\epsilon\in[0,1-1/e].\end{array}$
Proof. The logarithm function is concave, thus it follows that
$\log(1+\epsilon)\leq\log(1)+\epsilon\frac{d\log(x)}{dx}\bigg{|}_{x=1}=\epsilon.$
Moreover, by concavity of $\log(1-\epsilon)$ for $\epsilon\in[0,1-1/e]$, we have:
$\log(1-\epsilon)\geq\log(1)+\epsilon\frac{\log(1/e)-\log(1)}{1-1/e}=-\frac{\epsilon}{1-1/e}.$
Figure 13 gives a pictorial representation of the result.
Appendix B Additional details
B.1 Derivation of influence function
The influence function measures the sensitivity of a statistic to small changes in the data. The influence function of a statistic $T$ at a point $\zeta$ for a distribution ${\mathbb{P}}$ is given by:
$\text{IF}(\zeta)=\lim_{\varepsilon\to 0^{+}}\frac{T((1-\varepsilon){\mathbb{P}}+\varepsilon\delta_{\zeta})-T({\mathbb{P}})}{\varepsilon},$
where $\delta_{\zeta}$ is the Dirac distribution at the point $\zeta$. The influence function of entropic risk measure is given by:
@@ B.1 Derivation of influence function
$\begin{array}[]{lll}\text{IF}({{\boldsymbol{\xi}}})&=\lim_{\varepsilon\to 0^{+}}\frac{\log\left((1-\varepsilon){\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]+\varepsilon\exp(\alpha\ell({{\boldsymbol{\xi}}}))\right)-\log\left({\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]\right)}{\alpha\varepsilon}\\ &=\lim_{\varepsilon\to 0^{+}}\frac{\log\left((1-\varepsilon)+\varepsilon\frac{\exp(\alpha\ell({{\boldsymbol{\xi}}}))}{{\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]}\right)}{\alpha\varepsilon}=-\frac{1}{\alpha}\left(1-\frac{\exp(\alpha\ell({{\boldsymbol{\xi}}}))}{{\mathbb{E}}_{\mathbb{P}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]}\right).\end{array}$
Figure 14: Histogram of the $500$ loss samples, $\xi$, from a $\Gamma(10,0.24)$ distribution and the average IF of the entropic risk is computed by taking the mean of IF over scenarios in each bin with the darker blue shade representing higher (normalized) average IF.
When loss function $\ell(\xi):=\xi$ with $\xi\sim\Gamma(\kappa,\lambda)$, the influence function can be obtained in closed form as follows:
$\text{IF}(\hat{\xi})=-\frac{1}{\alpha}+\frac{\exp(\alpha\hat{\xi})}{\alpha\left(1-\lambda\alpha\right)^{-\kappa}}.$
Figure 14 presents the histogram of $500$ loss values generated from the $\Gamma(10,0.24)$ distribution. For each bin, we calculate the average influence function over samples and normalize these values to lie in the interval $[0,1]$, where darker shades represent higher average influence function. It can be seen that tail events have highest impact on the entropic risk but occur with very low probability. Therefore, these high-impact scenarios are likely not to be included in a finite sample, resulting in the underestimation of the entropic risk.
B.2 Bias mitigation using OIC
From Theorem 1 in Iyengar et al., (2023), it follows that for a loss function $h(t,\ell({{\boldsymbol{\xi}}}))$ with decision $t$: